OpenStudy (ammarah):

What term is needed to add to each side to complete the square? x^2+4x=5

OpenStudy (anonymous):

you subtract 4 from each side. because you can factor the expression \[x^2 +4x +4 \]

OpenStudy (anonymous):

into (x+2)^2

OpenStudy (ammarah):

how about 2x^2+4x=8

OpenStudy (ammarah):

@whpalmer4

OpenStudy (whpalmer4):

Well, you can factor out a 2 from everything, so you should do so.

OpenStudy (anonymous):

whoops im so sorry, I said subtract, but I meant add 4.

OpenStudy (ammarah):

i dont understand

OpenStudy (whpalmer4):

\[2x^2+4x=8\]Divide everything by 2 \[x^2+2x=4\] now complete the square

OpenStudy (ammarah):

how? subtract 4?

OpenStudy (whpalmer4):

Take half of the coefficient of \(x\). Square it. Add to both sides. Rewrite the left side as a perfect square \((x+a)^2\) where \(a\) is half of the coefficient of \(x\).

OpenStudy (ammarah):

so 6? how would i write it/

OpenStudy (whpalmer4):

what is 6?

OpenStudy (ammarah):

half of the coefficient is 1^ so 4+1=5

OpenStudy (whpalmer4):

okay, 5 isn't 6, but go on, 5 is better here :-)

OpenStudy (ammarah):

so the tern is 2 idk how to write it it just asks for the term

OpenStudy (ammarah):

is it (x+2)^2?

OpenStudy (whpalmer4):

Okay, here's how you check your work: multiply it out. Does it give you the right answer? If not, it's probably wrong. You haven't told me what your equation looks like after completing the square, so I can't do it for you.

OpenStudy (ammarah):

x^+2x=5

OpenStudy (whpalmer4):

Try typing that again.

OpenStudy (ammarah):

x^+2x-5?

OpenStudy (whpalmer4):

I have to go, so I'll just show you. \[x^2 + 2x = 4\]The coefficient of \(x = 2\), so \(a = 2/2 = 1\) We have to add \(a^2\) to both sides: \[x^2+2x+1 = 4+1\]\[x^2+2x+1 = 5\]Now you can rewrite the left side as \((x+a)^2\) \[(x+1)^2 = 5\] Check the work: \[(x+1)(x+1) = 5\]\[x^2 + 1x+1x + 1 = 5\]\[x^2 + 2x + 1 = 5\]\[x^2+2x +1-1 = 5-1\]\[x^2 + 2x = 4\] That matches our original equation after we factored out a \(2\) from all the terms, so it is correct.

OpenStudy (whpalmer4):

\[2(x^2+2x = 4)\]\[2x^2 + 4x = 8\]